OpenStudy (anonymous):

If a baseball player hits a baseball from 4 feet off the ground with an initial velocity of 64 feet per second, how long will it take the baseball to hit the ground? Use the equation h = -16t^2 + 64t + 4

OpenStudy (anonymous):

yeah like the ball really goes straight up and straight down. these problems are ridiculous

OpenStudy (anonymous):

when it hits the ground the height will be zero so set \[-16t^2+64t+4=0\] solve for t

OpenStudy (amistre64):

divide out the common of -4 to get: 4t^2 -16t -1 and lets work it :) 4(t^2-4t+4) = 1 + 4(4) (t-2)^2 = 17/4 t = 2 +- sqrt(17)/2 if i see it right t = 2 + sqrt(17)/2 should do it

OpenStudy (anonymous):

i made the mistake!

OpenStudy (anonymous):

should be \[4t^2-16t-1=0\] damn another wrong answer from me...

OpenStudy (amistre64):

its the simple stuff that gets me all the time :)

OpenStudy (anonymous):

which i have now deleted! i love revisionist history

OpenStudy (anonymous):

didn't get you, it got me. apparently i do not know how to divide 64 by 4 correctly

OpenStudy (anonymous):

Medal for "yeah like the ball really goes straight up and straight down. these problems are ridiculous"